OpenStudy (anonymous):

Write the equation of a line that has a slope of -4/3 and passes through the point (3, -2)

jigglypuff314 (jigglypuff314):

to find the equation of a line in slope-intercept form y = mx + b plug in what you know when m = slope = -4/3 and (x,y) = (3, -2) so y = mx + b (-2)= (-4/3)(3) + b find b then plug the number you found b to be back into the y = mx + b so y = (-4/3)x + (b)

OpenStudy (anonymous):

so y=-4/3x+6 is the answer

jigglypuff314 (jigglypuff314):

to find b (-2)= (-4/3)(3) + b multiply -2 = -4 + b then add 4 on both sides +4 +4 -2 + 4 = b

OpenStudy (anonymous):

what?

jigglypuff314 (jigglypuff314):